December 15, 2022

GOOD MUSIC AND BAD

A Lecture Given At Thomas Aquinas College

by MarcusBerquist

October 1991

There are two kinds of lectures. There are the lectures that you want to give, and the lectures that you’re talked into giving. This one is in the second category. It’s a good lesson to me to keep my mouth shut (something I should know at my age). I tend to speak vehemently on certain subjects, and find before long that I’m challenged to give a rational account of what I’ve just been asserting so passionately. Well, that’s how I got into this particular flx.

I’d like to start off with a little text from Plato, which doesn’t seem to have anything in particular to do with our subject, but I think we’ll see, as the discussion proceeds, that it has at least in general a great correspondence to our present discussion. Socrates is in jail, waiting for his execution, and some of his friends want to persuade him to escape, and go off to livein Thessaly and have a good time for the rest of his life. In the course of this conversation, his friend Crito says to him, ”You can see for yourself, Socrates, that one has to think of the opinion of the many as well. Your present position is quite enough to show that the capacity of ordinary people for causing trouble is not confined to petty annoyances, but has hardly any limits ifyou once get a bad name with them.” Socrates replies, ”I only wish that ordinary people had an unlimited capacity for doing harm, for then they might have an unlimited power for doing good, which would be a splendid thing, if it were so. Actually, they have neither. They cannot make a man wise or stupid, they simply act at random.”

I think that Socrates is stating a universal truth here that applies to many situations besides his own. One could state the truth this way: things which do us the greatest harm, are the very things which, when they are as they should be, do us the greatest good. Or, conversely, the things which do us the greatest good, are the very things which, when they are not as they should be, do us the greatest harm. Another way this truth is expressed is in the saying ”The corruption of the best is the worst”. Man is the noblest

of all the animals, but a bad man is much worse than a bad beast, one might even say, more beastly than any beast. Lilies that fester smell far worse than weeds.

The explanation for this, at least in general, is not far to seek. God arranged all things in number, weight, and measure, and he saw that it was exceedingly good. We see that the goodness of any created thing depends upon measure. When it becomes excessive or deficient with respect to that measure it becomes bad or evil. Thus the human virtues are all forms of moderation, for they deal with things whose goodness depends upon measure; for example, food. If so much food is good for you, it does not follow that more food is better. It’s possible to eat too much, and to eat too little, to eat when one shouldn’t, and not to eat when one should. The virtues moderate our desires in order to conform them to the measure of the objects themselves, a measure which, of course, must be discovered by right reason.

We can contrast our virtues in this respect with the theological virtues, which have God as their proper object; God, Who is infinite goodness. These virtues are not properly speaking means or middles. We cannot believe in God too much, we cannot hope in Him too much, we cannot love Him too much. But with respect to all created things, it is as we said. Their goodness depends on a kind of mean or moderation.

Tonight, we propose to speak about music, good music and bad music. We shall argue that there is good music, and explain as well as we can what makes the difference between good music and bad music. More particularly, we shall argue that good music at its best disposes the soul to virtue, moral and intellectual, while bad music does the opposite. I add that qualification ”at its best” for reasons that will be a little clearer later.

I would like to begin with two qualifications . First of all,we don’t say that good music produces virtue, but rather that it disposes one for it. Moral virtues, of course, are engendered by repeated actions of the same kind as those which come out of the virtue. One becomes brave by doing brave acts, and by doing

them for the right reason. Intellectual virtues are acquired by study and instruction. So ifmusic has a role to play in making men better, it’s a disposing role, rather than one which accomplishes that purpose. Secondly, we must note that there are other influences both for and against virtue that operate at the same time. There are other things that will dis pose your soul to virtue, or to vice, besides the music that you listen to. And this is both good news and bad news. Bad news, because bad music is only one of the many influences to vice. It very naturally goes along with the other influences, but it sometimes is in isolation from them for a given individual. But the effect of bad music may, for that individual, be mitigated by other influences that pull him in another direction. He will suffer loss because of his habit of listening. to such music, but perhaps not disaster. Further, we fmd that some individuals appear not to be affected very much at all by any kind of music. I’ve met people in my time who claim to like Mozart and Bach and Haydn quite well, but they also like jazz and rock music. And they tell me ”Italldepends on the mood I’m in.” I used to be very puzzled by that, and I still am, because it seems to me that the very thing about Mozart’s music that would delight them would cause them pain in this other music. There’s a saying ”He who adds a science, adds a pain.” The more perceptive you are about the good of things, the more you notice other things being out of harmony with that good, and suffer from it. Perhaps the truth here is that these individuals are not affected very much by anything, or at least by any music. As regards music, they are like shallow ponds inwhich you can have ripples but no waves. That was my first qualification.

My second qualillcation is with my method. In applying general considerations to particular cases we often need premises that are taken directly from perception. Perception cannot be communicated; either one has it, or one does not. Ifyoutell a man blind from birth that snow is white, you cannot communicate to him the perception that was the foundation of that statement. Either he has the sense of sight, which will reveal that to him, or not. But when perception is lacking, one can often argue from signs, from things which always or usually or naturally go with it, and may reveal it more clearly than it reveals itself, at least to a

given individual. For example, music often goes along with certain words, and certain kinds of dancing. It will sometimes happen that the words that go with the music and the dancing that spontaneously and naturally go with it will reveal to a given individual something in that music he might not otherwise have noticed. I read recently about a band or group which got in trouble with the law because the lyrics, ifyou can call them that, of their music were just too offensive to be listened to. But I heard no report that the music had been found too offensive to be listened to, and perhaps those who realized that the words were impossible should have realized that, after all, they went quite well with the music and revealed even more clearly than the music itself the actual character of what was being played.

We have to note also that custom has a certain effect on us here. When you hear something over and over again, it becomes customary, then ordinary, and then it seems normal. You get used to it. You half tune it out. Therefore people are often shocked when they hear the words or see the dancing which goes with music of a certain sort, because they’re not used to that. They’ve gotten used to the music, but the haven’t gotten used to the rest of it. So a lot of our more particular arguments tonight will be from signs, things that go with the good and the bad, and not so much from intrinsic definitions, or by giving a specific reasons in each case. Let us begin.

There are three places in the course of study where music is considered. The frrst, in the order of learning, is the liberal art of music, a part of the quadrivium, which examines music in the light of certain mathematical principles which it exhibits. Here, we see frrst in the order of learning that music is characterized by a reasonable order. We see in music a kind of unity and harmony between passions and reason. This kind of order is a good thing to see at the beginning because it’s proportioned to us. This is the easiest sort of order for us to see and appreciate. Even the young, without a great deal of experience, can apprehend an order of this kind.

The next place that music would come up in the course of study is in a way analogous to Aristotle’s consideration of tragedy in the Poetics,asa mode of imitation. This would be a thorough or definitive consideration in terms of the proximate genus, which is imitation, and the specific differences, imitation of what? and by what means? We can contrast this with the kind of treatment we fmd in the liberal art, wherea doctrine which is abstract and general is applied to a particular subject matter, a doctrine which has not been derived from that subject matter, and from consideration of its peculiarities, but from a more general and abstract consideration. The numerical ratios and proportions we study in harmony are common to music and other things as well.

The third consideration of music in the course of study is in ethics and political philosophy. We find this, for example, in Book VI of Plato’s Republic, in Book II of his Laws,andin Book VII of Aristotle’s Politics.Here, music is considered in terms of education. This is because, in the opinion of Aristotle and Plato and many others, music not only amuses and pleases, which is perhaps a sufficient reason for its being, as a kind of recreation and rest from life’s effortful activities, but it’s also dispositive. It has an effect on the soul for good or for ill. Therefore it pertains to ethics and politics to consider it. Itpertainsto education, which is concerned with the acquisition of virtue, which is of political and social as well as familial concern. We are all concerned that the citizens be good men.

Tonight, our consideration will be aimed at this third sort, the ethical significance of music, and the effect which the best sort of music has on character. However, we don’t argue, as we indicated before, that all good music is essentially of this sort. What pleases and refreshes, if it does not corrupt, is a good thing in moderation. Our life here is not like our life in beatitude, where we can always be doing the very highest thing. Here, we need a break. We need to sleep, and during our waking hours we need some recreation. We will not be maintaining that one ought to fmd some kind of moral significance in every piece of music that we enjoy. Perhaps it is there –I think in some way it is– but it is far more evident in one sort than in another. We do maintain that the best kind of music is

dispositive to virtue, and that music of the opposite sort tends to corrupt. The corruption of the best is the worst.

We can’t come immediately to this consideration. For we cannot sufficiently consider the end and ultimate good of anything unless we frrst consider its form, that is to say, its defining principles, and the power that comes from that form. Thus we must frrst consider music as a mode of imitation, and examine its specific power –how it affects us as music. The order, then, of our consideration, will be this. We will first of all examine the genus of music, which is imitation. Then we shall go on to consider its difference, especially as regards the object of imitation –what it is that music imitates. Then we shall consider the unique way in which music is like the object of imitation, and finally, the unique power it has by reason of that resemblance. The last consideration will lead to a consideration of the ethical or moral effect of music. We shall not, however, discuss the mathematical principles involved, except in a very limited way, and later on–the principal reason for this being my ignorance.

At the beginning of his Metaphysics,Aristotle divides art into three genera. There are the arts which produce necessary and useful things, for example carpentry. These are called ”servile” because they provide instruments of life and of the good life. Secondly, there are the arts which aim atrecreation and delight, what we now call the” fme arts”. The poet and the musician seek to please, and yet this pleasure is not the end of life. Life is not for the sake of recreation and amusement. Thirdly, there are the arts whose end is knowledge, for example, geometry. The proper end of geometry is to know about magnitude and figure. Such arts are called ”liberal” because they befit a free man, having value in themselves, because in and of themselves they make us know something worth knowing. And knowing is, largely, the end of life.

The difference between the frrst and the third of these genera is evident. In the frrst genus, knowledge is simply for the sake of the making. If we could have the product without the knowledge we would not bother about the knowledge. But in the third case the end sought is the knowledge itself. But as regards the second

genus, the fine arts, the contrast is not so clear. Middles, things in between, are always hard to define. But we can say this. The poet, the painter, and the musician are like the carpenter in this respect at least: their knowledge is for the sake of some work. The art of the poet is for the sake of the poem that he composes or the play or the story, the art of the musician is for the sake of the composition that he makes, the art of the sculptor is for the sake of the statue, and so on. The end of these sciences is making, not knowing, and making is for the sake of the thing made. Nevertheless, when we consider the use of the products of these arts, the fine arts, we perceive a certain likeness to the liberal arts. For the use of the products is not ”use” in the ordinary sense; it’s in being seen or being heard, that is to say, in some act of knowledge, and knowing something is not using it in the ordinary sense. Thus, when we listen to the poet or the musician, we are not using his product to bring about some further effect by means of it, but we are ourselves being affected. But this affecting is in the frrst instance a kind of knowing, involving both sense and intellect. Thus we see why these arts are called ”fme,” at least when compared with the servile arts. The use of the products, here, is a sort of knowing.

Next, we ask what do these fine arts produce such that they should receive such a use? Aristotle, at the beginning of the Poetics, gives an answer. He says this: ”Our subject being poetry, I propose to speak not only of the art in general, but also its species and their respective capacities”. Then he goes on a little bit further: ”Epic poetry and tragedy, as also comedy, dithyrambic poetry, and most flute-playing and lyre-playing, are all, viewed as a whole, modes of imitation. But at the same time they differ from one another in three ways: either by a difference of kind in their means, or by differences in their objects, or in the manner of their imitation”. (Poetics 1447a7-18)

From the discussion that follows in Aristotle’s text, we see that Aristotle also regards painting and sculpture as modes of imitation. Aristotle, then, regards this as the most general and fundamental difference between the fine arts and the other arts. Accordingly, an artist of this sort is essentially a maker of imitations. And thus Aristotle goes on to differentiate among the fine arts on this basis.

These arts differ from one another by a difference in the objects imitated, the means of imitation, and the manner of imitation.

This view of the fme arts, that they are arts of imitation, is quite probable on the face of it, and seems to fit the prominent examples which are taken to be typical. It has been the general view throughout most of our history. But there are objections and difficulties, and they are more acutely urged in modern times. Let me state a few of them.

First of all, one says, not all art is representative. So the difference is not universal. We see that those artists that can and do produce representations also produce non-representational art. So the difference is a dividing difference rather than a constituting difference.

Secondly, one argues, when we consider even representational art, it seems that the artist has not made a reflection or reproduction of his original, but has transformed it. His work is significant as a work of art not by what he has taken or received from the original, but by what he has originated himself in the product. Anexample can be taken from photography and painting. Almost everyone regards painting as more of an art than photography, and yet if we consider just the literal and exact likeness of the object being portrayed it seems that photography is more exact than painting. It seems, on this view, that the likeness of the imitation to its original is less significant than the difference. Along the same lines, allseem to recognize that in some way art improves upon nature, that the imitation improves upon the original. Thus we tend to see the artist as creative, even with his representational work. Accordingly, although not all would draw this conclusion, creationrather than representation seems to be the perfection of artistic activity. The work originates in the artist’s personal vision.

This view, however, cannot be sustained. We wonder, to begin with, why one says”create” and not ”make”. The choice of that word is, I think, significant. The terms ”create” and ”creation” are terms which have a peculiar weight and significance in the Western

tradition, because of the Catholic faith and sacred doctrine. To apply these terms to the artist, the maker, is to attribute to him, in one way or another, and more or less, what is uniquely true of God’s efficient causality. That is to say, to call the artist ”creator” is to consider him Godlike in that respect in which he had previously been considered most unlike God. For God’s making of the world is called ”creation” specifically for two reasons. First of all, He does not make by transforming any pre-existent matter. That’s what’s usually meant by ”making the world out of nothing”. Secondly, He does not work from any exemplar, or ideal, or model other than Himself. That is, the form by which he acts is simply His own form, in no way acquired from another. Thus creation, properly understood, means a making in which there is no principle other than the maker himself.

Now our making, by contrast, requires both a matter and a form, neither of which we originate, but only find. This is obviously true as regards the matter. Not even the most enthusiastic partisan of artistic creativity maintains that the artist can make anything out of no material. But as regards the form, it might seem that there is a kind of true originality, and that with respect to form that the artist is indeed a creator. But this can’t be so. For the human intellect is natively a blank. It has no forms of itself by which it might know and act. This is why Aristotle compares it to prime matter; it is purely potential in its own order. It is that which can be understanding, but is not. It becomes actually understanding by what it receives from experience – certain forms and notions that belong to the objects of experience and are not originated by the artist himself. Therefore, since the maker